0.06x+0.09(x+200)=63
Answer by umail08(87)

0.06x + 0.09(x+200) = 63
0.06x + 0.09x + 18 = 63
0.15x + 18 = 63
0.15x + 18 – 18 = 63 -18
0.15x = 45
0.15x/0.15 = 45/0.15
x = 300